# JEL: C81, O32

### v1

@@ 5. Discussion
Existing work documents patterns that are not inconsistent with those in this paper. However, we offer an alternative interpretation. Like Park et al., (2023) and Bloom et al., (2020), we find a substantial increase in the size of the medical research ecosystem. Park et al., (2023) document that this increase in scale brings with it a decline in the average importance of published scientific articles. Technically, we find the same. As increases in non-clinical trials outpace (non-existent) changes in clinical trials, the average productivity of studies in Figure 3 mechanically decreases. In debates on productivity, however, we argue that this measure of central tendency is not the policy-relevant object. To be sure, these findings raise a host of other important questions—about misaligned incentives, wasteful spending, and the productive efficiency of science. But canonical frameworks, such as Romer, (1990), center on the simple idea that it is high-intensity investments that generate meaningful shifts in outcomes. An explosion of low-capital research that has a negligible impact on outcomes, then, is not cause for concern from a growth perspective. That is, growth-oriented studies may prefer to focus—like much of the research on science and innovation—on the right-tail of the productivity distribution (e.g., Kelly et al.,, 2021).
@@
Our findings do not speak to shifts in upstream, basic science inputs into clinical trials, or to shifts in factor prices. In our view, there is little credible evidence indicating changes on either dimension during this time period. This setting allows us, only, to note that any shift in inputs or factor prices must be perfectly offset by a countervailing force, to account for the complete stability of clinical trial counts, citations, journal placements, and reliance on public funding in our data.
Our findings, also, apply to roughly one decade: 2010 to 2022. We impose this restriction because policy changes in the early 2000s substantially shifted incentives for both publication and standardized formatting of clinical trial results. See Laine et al., (2007) and Schulz et al., (2010) for details. Given our methods, these changes mean that data construction over a longer period risks decreasing the quality of our data. Thus, we cannot speak to trends that precede this period. That clinical trial investment is remarkably stable in the past decade, however, is likely the relevant consideration for contemporary public policy.
@@
What our findings, and the simple framework borrowed from the growth literature, do suggest is that there is likely value in efforts to increase the productivity and decrease the costs of clinical trials. A long-standing discussion in medicine and statistics, spurred by Altman, (1994), centers on the argument that “we need less research, better research, and research done for the right reasons.” Our data indicate that nearly half of all clinical trials published in our sample are never cited by leading medical journals, and roughly 15 percent are never cited at all. This heterogeneity is consistent with findings in Ioannidis, 2016b and van Zwet et al., (2023), who imply that there may exist trade-offs between quantity and data quality in this setting. If increasing the scale of research infrastructure is a policy objective, such efforts should invest in tools, like patient registries, that enable the production of high-quality, representative information (Alsan et al.,, 2024; Durvasula,, 2023).
@@
Supplemental Appendix to: Medical Research as a Productivity Indicator [^1] Date: May 12, 2024
@@
Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bsubsection.B.1subsection.B.2subsection.B.3subsection.B.4subsection.B.5subsection.B.6appendix.Cappendix.D
@@ Appendix A Publication Data
In this Appendix, we detail our treatment of the PubMed / MEDLINE database (hereafter, “PubMed”). We give a general overview of the PubMed data in Section A.1. In Section A.2, we describe the process of identifying our initial sample of abstracts, i.e., the “Potential Universe” displayed in Figure 4. We discuss the process of determining the set of leading medical journals in Section A.3.
@@ A.2.5. Novel Census
We add columns to Tables A.2, A.3, A.1, and A.4 that document the frequency of each flag in the data constructed for this paper (we use the conservative sample). Several facts are worth noting. First, Table A.2 suggests that—between 2010 and 2022—less than 25 percent of records identified as reporting the results of a clinical trial reported a registry identifier in their abstract. Although, in principle, registry identifiers may be reported in publication full-texts, rather than abstracts, the low frequency of identifiers in our sample suggests non-compliance with requirements, consistent with patterns in ClinicalTrials.gov documented in DeVito et al., (2020). Second, Table A.1 sheds light on the differences in trial composition that we observe in Figure 1: NLM tags that, ostensibly, capture clinical trials have little overlap with one another and do not fully capture the records in our sample. Of the roughly 150,000 records in our census, approximately 100,000 could be identified using NLM tags for “clinical trial” (and its variants) and “randomized controlled trial.” Observe, however, that use of such flags captures a much larger number of records than one would intend to include. Finally, Table A.4 suggests that although various text-based flags that may indicate a record is a clinical trial, in the sense of 3.1, are correlated, there is less overlap than one might expect.
@@ A.3. Determining the Set of Leading Medical Journals
Angrist et al., (2020) propose a strategy to identify leading journals in a field, which we adapt to this context. In medicine, the “trunk journals,” per Angrist et al., (2020), are the Journal of the American Medical Association and the New England Journal of Medicine. We collect all citations originating from these journals to records in PubMed between 2010 and 2022. Angrist et al., (2020) collect initial journal lists for their citation exercise of the journal is one of the top fifty most cited by a trunk journal. As medicine includes a large number of subfields, we instead collect a list of journals that receive at least 100 citations from a trunk journal over this time period, which yields a list of 84 journals.
@@ A.4. The Web of Science
For each paper in our sample, we extract countries from available mailing addresses for first- and last-listed authors. First- and last-authors have the same listed country for 85 percent of clinical trial records and 91 percent of non-clinical trial records. Thus, we impute to each record the country associated with the first-listed author. In Figure 6, we plot geographic trends over time.
Appendix B Prompt Design, Fine-Tuning, and Performance Assessment
@@
In this appendix, we detail the construction of the sample of clinical trials studied in Section 4. We begin in Section B.1 by describing the process we used to hand-label a sample of publications according to whether they satisfy the restrictions enumerated in 3.1. In Section B.2, we document poor performance of several approaches for classifying publications according to whether they belong in our sample based on standard machine learning methods. Section B.3 overviews the process of constructing prompts that perform well when used to query proprietary language models. Details associated with fine-tuning an open-source language model are given in Section B.4. Additional measurements of the performance of our model are given in Section B.5.
B.1. Hand-Labeling
@@
We devise an interface for abstract labeling, which allows us to review records quickly. Figure B.1 provides an example of this interface. A labeller reads in a set of publication records and is shown one PubMed identifier (PMID) and one abstract at a time. The labeller can select that the abstract satisfies our sample criteria, or else can indicate a reason that the abstract should be excluded. The buttons for exclusion mirror the inclusion and exclusion criteria in 3.1, using short-hand convenient for labellers. We exclude records that report, only, the protocol for a clinical trial without results (“protocol”), that report the results of an observational (“observational”) or retrospective (“retrospective”) study, that report the results of a study involving animals (“animal”), that do not include human subjects (“no human subjects”), [^36] In practice, “animal” and “no human subjects” catch different sets of records. We flag records that include any animal with “animal.” Many studies in our sample of candidate records occur in vitro—in test tubes or other laboratory-based settings. There, no living subjects of any kind are enrolled and the “no human subjects” button is used. that do not study a drug (“no drug”), that report the results of a specific clinical case or patient experience (“case report”), or that summarize the findings of existing work (“meta-analysis”). By our own estimates, the labelling interface increased the speed of labelling by a factor of seven. Each record was reviewed, and labelled, twice.
@@
Figure B.1. Abstract Labeling Interface
Notes: Figure B.1 displays a screenshot of the application we developed to hand-label abstracts. Each abstract was labelled according to whether it satisfied the sample criteria enumerated in 3.1. If a publication did not satisfy our sample restrictions, we stipulated a reason. The user interface was developed with the “Meerkat” Python package available at the link: https://github.com/HazyResearch/meerkat.
B.2. Benchmark Comparison to Standard Machine Learning Methods

### v2

@@ 5. Discussion
Existing work documents patterns that are not inconsistent with those in this paper. We differ, however, in interpretation. Like Park et al., (2023) and Bloom et al., (2020), we find a substantial increase in the size of the medical research ecosystem. Park et al., (2023) document that this increase in scale brings with it a decline in the average importance of published scientific articles. Technically, we find the same. As increases in the quantity of non-trial publications outpace (non-existent) changes in clinical trials, the average productivity of studies in Figure 3 mechanically decreases. We find, also, that increases in the quantity of non-trials are coincident with declines in various measures of publication quality. From the perspective of a productivity calculation, however, we argue that measures of average quality provide little insight. We cannot rule out, for example, that the same amount of research is being produced in these non-trial publications, but that what had previously been reported in a single paper is now spread across several distinct records. To be sure, these findings raise a host of other important questions about scientific research—about misaligned incentives, wasteful spending, and both productive and allocative efficiency. But on the issue of research productivity, these shifts provide very little insight.
@@
Our findings do not speak to changes in upstream, basic science inputs into clinical trials, or to changes in factor prices. As Section 3 discusses, in our view, there is very little credible evidence that might allow us to investigate changes on either dimension during this time period. We can note, only, that any shift in inputs or factor prices must be perfectly offset by a countervailing force, to account for the complete stability of clinical trial counts, citations, journal placements, and reliance on public funding in our data.
Our findings, also, apply to roughly one decade. We impose this restriction because a set of policy changes in the early 2000s substantially shifted incentives for disclosure and publication of clinical trial results (Laine et al.,, 2007; Schulz et al.,, 2010). To be clear, there is nothing about our method that limits its applicability to data reported in the particular form required by these policy changes. Our concern here is about the interpretability of the resulting series. As noted previously, proxies for scientific research provide very little insight if the relationship between one unit of the proxy and units of underlying, unobserved research shift over time. Policy changes in this sector are such that we risk introducing this source of bias with a longer series. Thus, we cannot speak to trends that precede this period. From the perspective of contemporary public policy, however, trends from 2010 to the present are likely to be most relevant.
We reconcile two stylized facts about research and inventive output in a manner that is both optimistic and pessimistic. On the one hand, we find that, in the most recent period, both the inputs and outputs of an “idea production function” (2.1) are constant across periods. This should alleviate the concern, in this sector, that “ideas are getting harder to find” (Bloom et al.,, 2020).
On the other hand, it is worth remarking, briefly, on the implications of completely stable clinical trial trends. The early twenty-first century is often described by scientists as a “golden age” for medicine. Breakthroughs, such as the completion of the Human Genome Project and the diffusion of DNA sequencing technologies one billion times faster than their predecessors, rapidly expanded the set of available scientific opportunities (e.g., Scannell et al.,, 2012). Technological advances were matched by institutional support, including the doubling of the budget of the U.S. National Institutes of Health between 1998 and 2003. Against these factors that seemingly forecast a period of rapid innovation, the two facts documented in this paper—stability in productivity-relevant research investments, but rising quantities of comparatively low-impact forms of research—highlight the importance of understanding frictions that affect the translation of basic scientific insights into medical technologies relevant to patients.
@@
Specifically, these patterns suggest that there is likely value in efforts to increase the productivity and decrease the costs of clinical trials, and in efforts to reconcile incentives in academic publishing with social value. A long-standing discussion in medicine and statistics, spurred by Altman, (1994), centers on the argument that “we need less research, better research, and research done for the right reasons.” Our data indicate that nearly half of all clinical trials published in our sample are never cited by leading medical journals, and roughly 15 percent are never cited at all. Outside of clinical trials, our findings suggest that the quantity of low-quality research might be growing over time. This heterogeneity is consistent with findings in Ioannidis, 2016b and van Zwet et al., (2023), who imply that there may exist trade-offs between quantity and data quality in this setting.
If increasing the scale of research infrastructure is a policy objective, such efforts should invest in tools, like patient registries, that enable the production of high-quality, representative information (Alsan et al.,, 2024; Durvasula,, 2023).
@@
Supplemental Appendix to: Medical Research as a Productivity Indicator [^1] Date: August 3, 2024
@@
Contents  appendix.Asubsection.A.1subsection.A.2subsubsection.A.2.1subsubsection.A.2.2subsubsection.A.2.3subsubsection.A.2.4subsubsection.A.2.5subsection.A.3subsection.A.4appendix.Bappendix.Csubsection.C.1subsection.C.2subsection.C.3subsection.C.4subsection.C.5subsection.C.6appendix.Dsubsection.D.1subsection.D.2subsection.D.3subsubsection.D.3.1subsubsection.D.3.2subsubsection.D.3.3appendix.E
@@ Appendix A Publication Data
In this Appendix, we detail our treatment of the PubMed / MEDLINE database (hereafter, “PubMed”). We give a general overview of the PubMed data in Section A.1. In Section A.2, we describe the process of identifying our initial sample of abstracts, i.e., the “Potential Universe” displayed in Figure 4.1. We discuss the process of determining the set of leading medical journals in Section A.3.
@@ A.2.5. Novel Census
We add columns to Tables A.2, A.3, A.1, and A.4 that document the frequency of each flag in the data constructed for this paper (we use the conservative sample). Several facts are worth noting. First, Table A.2 suggests that—between 2010 and 2022—less than 25 percent of records identified as reporting the results of a clinical trial reported a registry identifier in their abstract. Although, in principle, registry identifiers may be reported in publication full-texts, rather than abstracts, the low frequency of identifiers in our sample suggests non-compliance with requirements, consistent with patterns in ClinicalTrials.gov documented in DeVito et al., (2020). Second, Table A.1 sheds light on the differences in trial composition that we observe in Figure 3.1: NLM tags that, ostensibly, capture clinical trials have little overlap with one another and do not fully capture the records in our sample. Of the roughly 150,000 records in our census, approximately 100,000 could be identified using NLM tags for “clinical trial” (and its variants) and “randomized controlled trial.” Observe, however, that use of such flags captures a much larger number of records than one would intend to include. Finally, Table A.4 suggests that although various text-based flags that may indicate a record is a clinical trial, in the sense of 3.1, are correlated, there is less overlap than one might expect.
@@ A.3. Determining the Set of Leading Medical Journals
Angrist et al., (2020) propose a strategy to identify the leading journals in a field, which we adapt to this context. In medicine, the “trunk journals,” per Angrist et al., (2020), are the Journal of the American Medical Association and the New England Journal of Medicine. We collect all instances in which papers published in these journals cite other records in PubMed, between 2010 and 2022. In Angrist et al., (2020), the authors select the fifty journals receiving the most cites from these trunk journals. As medicine contains a large number of subfields, we instead collect a list of journals that receive at least 100 citations from a trunk journal over this time period, which yields a list of 84 journals.
In our data, papers published in trunk journals cite records in 3,136 unique journals. [^61] We use unique entries in the field medlineta—which correspond to distinct MEDLINE abbreviations for journal names—to identify distinct journals. See https://www.nlm.nih.gov/bsd/serfile_addedinfo.html for details on name abbreviations in this database. We confirm that the count of journals is identical if we, instead, use journals identified by the National Library of Medicine’s unique journal identification number. The median journal is cited twice by these two trunk journals over our time period. The most frequently cited journals are, perhaps unsurprisingly, the trunk journals themselves: JAMA receives 3,143 citations, and the New England Journal of Medicine receives 4,302 citations. The next ten journals, by frequency of citations, are: Lancet (1,457), Circulation (789), Journal of Clinical Oncology (734), Nature (660), Blood (651), PLoS One (645), Journal of the American College of Cardiology (593), Clinical Infectious Diseases (589), BMJ (586), and Annals of Internal Medicine (586).
@@ A.4. The Web of Science
For each paper in our sample, we extract countries from available mailing addresses for first- and last-listed authors. First- and last-authors have the same listed country for 85 percent of clinical trial records and 91 percent of non-clinical trial records. Thus, we impute to each record the country associated with the first-listed author. In Figure 4.4, we plot geographic trends over time.
We use the Web of Science to collect details about author career and publication histories, which allow us to investigate patterns in the production of medical research. The Web of Science assigns to authors a unique, persistent identifier (DAIS ID), which allows us to collect information on authors’ output before and after the publication focal records in our sample. To our knowledge, there have been no large-scale efforts to validate the quality of these researcher identifiers. On inspection, it appears as though the identifiers perform best, perhaps unsurprisingly, for researchers with distinctive names. When researchers have names that are more frequent in the population, our spot-checking (against online publication records) suggests that there are more likely to be errors. Nonetheless, tracking researchers using these identifiers is standard in this literature, as there are no obviously superior alternatives. Especially given that researcher productivity over time is not a centerpiece of our analysis, we do not undertake additional validation exercises.
Appendix B ClinicalTrials.gov
ClinicalTrials.gov is a registry of clinical trials, run by the United States National Library of Medicine. As of 1 August 2024, ClinicalTrials.gov lists more than 500,000 studies in all 50 states and in 222 countries. [^63] See https://clinicaltrials.gov/about-site/trends-charts. ClinicalTrials.gov was available for study registration beginning 29 February 2000. At this time, there were 1,255 registered studies. [^64] In 1988, in response to public pressure surrounding perceived lags in the development of therapeutics and preventives to treat HIV/AIDS, U.S. Congress passed the Health Omnibus Programs Extension Act (Public Law 100-607), which required the creation of a database of AIDS Clinical Trials Information Services (ACTIS). Publicly-funded clinical trials that were associated with these previous data registration efforts were included in ClinicalTrials.gov at its inception. By the beginning of 2006, there were 24,822 studies and by the beginning of 2012, 118,020. In January 2024, there were 477,227. The geographic composition of these studies has shifted considerably over time, as have reporting patterns from various firms and universities.
In this Appendix, we use data from ClinicalTrials.gov to illustrate its limitations as a source for the type of high-quality records necessary for studies of productivity. Given the parallels between this paper and a long literature that studies the usefulness of patent data for studies of innovation, it is worth highlighting the important, perhaps obvious, distinction between ClinicalTrials.gov as a government database and records collected by the U.S. Patent and Trademark Office (USPTO). Records of U.S. patents are collected by the USPTO itself. Disclosure is widely considered a pivotal component of the patent system: in exchange for a short-term quasi-monopoly, information on patented inventions is placed in the public domain. Although ClinicalTrials.gov is described in similarly forceful terms in the laws that authorize its creation, disclosure has been—and continues to remain—effectively voluntary, in the sense that legal tools have never been used by a federal agency to compel the registration of a study or disclosure of its findings.
Institutional details suggest that the “voluntariness” of this reporting has shifted over time, in ways that likely affected sponsors’ propensity to register. [^65] In guidance to researchers on the use of ClinicalTrials.gov, Tse et al., (2018) collect a list of the ten challenges associated with using these data, chief among which is the fact that changes in medical journal reporting requirements and federal law make it extremely difficult to determine if increases in the size of the database are related to actual increases in the amount of scientific research. A coalition of medical journal editors began encouraging registration in 2005 (Laine et al.,, 2007), registration requirements were imposed for certain FDA-regulated trials in 2008, and the FDA began threatening to impose fines on noncompliant sponsors in 2021. In Table B.1, we collect a list of five changes in the past 25 years that changed sponsors’ requirements for trial registration.
Table B.1. Policy Changes Affecting the Composition of ClinicalTrial.gov
Policy (Effective Date) Registration Requirement
Food and Drugs Administration Modernization Act (21 November 1997; ClinicalTrials.gov registry available starting on 29 February 2000) Clinical trials of investigational new drugs for serious or life threatening conditions and diseases
International Committee of Medical Journal Editors (1 July 2005 for newly initiated trials and 13 September 2005 for ongoing trials) All clinical trials as a condition for consideration of publication of results
Food and Drugs Administration Amendments Act (FDAAA) (27 September 2007) Non-phase 1 clinical trials of FDA regulated drug and biological products, and non-feasibility trials of FDA regulated device products
Final rule implementing FDAAA in Title 42, Part 11 of the Code of Federal Regulations (42 CFR Part 11) (18 January 2017) Non-phase 1 clinical trials of FDA regulated drug and biological products, and non-feasibility trials of FDA regulated device products
Notes: Table B.1 collects a list of five policy changes in the past 25 years that changed sponsors’ requirements for trial registration. This table is adapted from Tse et al., (2018).
We inspect a version of the Aggregate Analysis of ClinicalTrials.gov (AACT) data, a publicly available relational database that contains information on all studies registered on ClinicalTrials.gov, downloaded on 2023 May 31. Of interest, for us, is the question of whether ClinicalTrials.gov can provide an accurate count of the total quantity of clinical trials and, if so, at what date such an accurate count might be available. Figure B.1 displays the results of our investigation.
Figure B.1. Number of Entries in Clinical Trials.gov, by Sample Restrictions
Notes: Figure D.1 displays the counts of the number of records in ClinicalTrials.gov that satisfy various sample restrictions. The $x$-axis of the top panel is the year in which the clinical trial was completed. The $x$-axis of the bottom panel is the year in which the clinical trial was registered on ClinicalTrials.gov.
We plot counts of records, selected in various ways, against two x-axes. In the top panel of Figure B.1, we plot counts associated with the year in which clinical trials are marked complete. [^66] These completion dates are somewhat imprecise. If a trial is registered, but no results are reported, the completion date is an estimate, based on the projected completion date in the study’s protocol. If the trial is registered and has associated results, the completion date is the actual date on which data collection ended. In the bottom panel, we plot counts based on the year in which the study was posted on ClinicalTrials.gov. Measures of trial completion are informative, insofar as they capture the date on which information production is finished. However, trials vary greatly in length—the key observation in Budish et al., (2015)—and as a result represent investments “sunk” at very different earlier points in time. Measures of trial posting are analogous, in some ways, to publication, in that they capture decisions to disclose information. Even without considering the specific series in Figure B.1, observe that the bottom panel includes a variety of sharp changes in the count of records. These correspond to several of the policy changes in Table B.1. We see a sharp change in all four series in 2005, when medical journal policies were changed, a sharp change in 2008, when the FDAAA took effect, and a sharp (albeit smaller) change in 2016, in the lead-up to the effective dates of the final two policy changes in Table B.1. We observe a smaller, but discernible, change in the slope of each series in 2008 when measures of trial completion are used. We draw attention to these differences, in part, because an important challenge for the use of these records in research is determining the appropriate measure of time.
Turning to the content of these series, we begin by examining the total number of records indexed in ClinicalTrials.gov over time (light green). We define a record as a unique “national clinical trial identifier,” although Tse et al., (2018) note that one study may be assigned multiple NCTs. There is no method, to our knowledge, that further distinguishes between records. In both panels, this series begins at zero in the year 2000 and increases sharply, to roughly 30,000 records by the end of our data, 2022. On inspection, it is clear that many of the 30,000 records included in the 2022 bin are not clinical trials. Among those that are clinical trials, many do not study medicines, but rather randomize exposure to interventions involving psychological treatments and exercise.
We impose three sets of restrictions, which aim to bring the data closer to a consistent sample of clinical trials studying the effects of a drug. We select these restrictions based on examination of various fields in ClinicalTrials.gov and with the aim of providing an illustration of our concerns with the use of this dataset “off the shelf.” There are alternative ways of constructing a sample of records that may be necessary to yield relevant data for different research questions.
We begin by dropping a set of records for which the study “status” variable in the database suggests that the study was either never started or was terminated. Specifically, we drop 40,622 records for which the status is listed as “Withdrawn,” “Suspended,” or “Terminated.” This status restriction is plotted in teal and has little impact on the trends in either panel.
Next, we observe that not all records in ClinicalTrials.gov are interventional studies. We drop any study with a type other than “interventional.” This removes an additional 98,138 observations from our data. This status and type restriction is plotted in blue. We observe that the sharply increasing trend in the later years of our data flattens out slightly. That is—the growth in the quantity of records is driven, in part, by growth in the number of registered non-interventional studies.
We implement a third restriction. We observe that clinical trial “phases” are a hallmark of drug trials: medical devices, for example, are approved via a “stage” process. Any record for which the trial phase is either missing or “Not Applicable” is then less likely to be a drug. We drop these records. This removes 151,541 observations from our data. We plot the resulting series in purple. When all three restrictions are imposed, we find steady increases in the quantity of trials leading up to 2008 and a stable series from 2008 forward. Reassuringly, this trend mirrors that documented in Figure 4.1. These plots suggest a level difference of roughly 2,000 observations, where Figure 4.1 is higher. This is consistent with the idea that ClinicalTrials.gov is not a global registry and, as Zarin et al., (2017) notes, many sponsors continue not to register studies in ClinicalTrials.gov
We emphasize that it is not obvious—in the absence of corroborating evidence, similar to that in Figure 4.1—that the version of this series with all three restrictions in Figure B.1 provides an accurate measure of trends in research, especially given the sharp changes in the early part of the series and the somewhat haphazard way that we identified a set of sample restrictions.
A series of recent papers—including DeVito et al., (2020) and DeVito and Goldacre, (2021)—document limitations with the quality of particular data elements, which further suggests the difficulty of using this database as a standalone source of estimates.
Appendix C Prompt Design, Fine-Tuning, and Performance Assessment
@@
In this appendix, we detail the construction of the sample of clinical trials studied in Section 4. We begin in Section C.1 by describing the process we used to hand-label a sample of publications according to whether they satisfy the restrictions enumerated in 3.1. In Section C.2, we document poor performance of several approaches for classifying publications according to whether they belong in our sample based on standard machine learning methods. Section C.3 overviews the process of constructing prompts that perform well when used to query proprietary language models. Details associated with fine-tuning an open-source language model are given in Section C.4. Additional measurements of the performance of our model are given in Section C.5.
C.1. Hand-Labeling
@@
We devise an interface for abstract labeling, which allows us to review records quickly. Figure C.1 provides an example of this interface. A labeller reads in a set of publication records and is shown one PubMed identifier (PMID) and one abstract at a time. The labeller can select that the abstract satisfies our sample criteria, or else can indicate a reason that the abstract should be excluded. The buttons for exclusion mirror the inclusion and exclusion criteria in 3.1, using short-hand convenient for labellers. We exclude records that report, only, the protocol for a clinical trial without results (“protocol”), that report the results of an observational (“observational”) or retrospective (“retrospective”) study, that report the results of a study involving animals (“animal”), that do not include human subjects (“no human subjects”), [^67] In practice, “animal” and “no human subjects” catch different sets of records. We flag records that include any animal with “animal.” Many studies in our sample of candidate records occur in vitro—in test tubes or other laboratory-based settings. There, no living subjects of any kind are enrolled and the “no human subjects” button is used. that do not study a drug (“no drug”), that report the results of a specific clinical case or patient experience (“case report”), or that summarize the findings of existing work (“meta-analysis”). By our own estimates, the labelling interface increased the speed of labelling by a factor of seven. Each record was reviewed, and labelled, twice.
@@
Figure C.1. Abstract Labeling Interface
Notes: Figure C.1 displays a screenshot of the application we developed to hand-label abstracts. Each abstract was labelled according to whether it satisfied the sample criteria enumerated in 3.1. If a publication did not satisfy our sample restrictions, we stipulated a reason. The user interface was developed with the “Meerkat” Python package available at the link: https://github.com/HazyResearch/meerkat.
C.2. Benchmark Comparison to Standard Machine Learning Methods